OpenStudy (anonymous):

Change the repeating decimal 0.242424242424 . . . to a reduced fraction.

OpenStudy (anonymous):

thats an irrational number

Directrix (directrix):

Let x = .242424... Then 100x = 24.24.24... x = .242424... ---> The object is to multiply by a power of ten that will line up the repeating blocks and subtract them off. ------------------------------- 99x = 24 x = 24/99 x = 8 / 33

OpenStudy (anonymous):

thanks ;))